OpenStudy (anonymous):

What is the value of n in the equation five-sixths equals n divided by five plus one-half

OpenStudy (anonymous):

so the equation is : 5/6= n/5 + 1/2?

OpenStudy (campbell_st):

subtract 1/2 from both sides of the equation and you get 2/6 = n/5 multiply both sides of the equation by 5 to get n

OpenStudy (anonymous):

yes @Jaodao

OpenStudy (anonymous):

5/6-1/2=n/5 5-3/6=n/5 2/6=n/5 1/3=n/5 5/3=n therefore, n = 5/3

OpenStudy (anonymous):

isnt the answer a whole number

OpenStudy (anonymous):

not with the following equation its not

OpenStudy (anonymous):

but the answer u got is not on the choices ii got

OpenStudy (anonymous):

could you perhaps write the equation once yourself? that'd would help

OpenStudy (anonymous):

5/6=n/5+1/2 i have 3/5 on one of the choices

OpenStudy (anonymous):

im getting n=10/6=5/3 which is the same result that @campbell_st got. for the answer to be 3/5, n would be represented as 1/n but that is not the case.

OpenStudy (anonymous):

can u please xplain it to me

OpenStudy (anonymous):

5/6 = n/5+1/2 is the equations and we need to find n subtract 1/2 from both sides 5/6-1/2= n/5 +1/2 -1/2 to get: 5/6-1/2= n/5 now find multiply (-1/2 x 3/3) to get (-3/6) we do this to make the denominator the same value as the other number, so... 5-3/6=n/5 we get 2/6=n/5, now multiply both sides with 5 so that we eliminate denominator on 'n' 2x5/6 = nx5/5 we gives 10/6=n therefore by simplifying 10/6 we get n = 5/3 that is the logic behind it, if this is not the answer then there must be a mistake with the equation or the question or answer or me and campbell